I need help to solve this problem
Two ships leave port at the same time, one heading due south and the other heading due east. Several hours later, they are 170 miles apart. If the ship traveling south traveled 70 miles farther than the other ship, how many miles did they each travel?

Answer by stanbon(48552)

Two ships leave port at the same time, one heading due south and the other heading due east. Several hours later, they are 170 miles apart. If the ship traveling south traveled 70 miles farther than the other ship, how many miles did they each travel?
--------------
Draw the picture.
You have a right triangle with the south leg being x+7 and the east leg
being x; the hypotenuse is 170.
---------
Use Pythagoras to get:
x^2+(x+70)^2= 170^2
x^2+x^2+140x+4900 = 170^2
2x%2+140x+-24000 = 0
x^2+70x-12000=0
(x+150)(x-80)=0
Positive answer is x=80 miles
x+70 = 150 miles
==============
Cheers,
Stan H.